A 2-kg mass of oxygen expands at a constant pressure of 172 kPa in a piston-cylinder system from a temperature of 32 C to a final temperature of 18 C. Determine (a) the heat required; (b) the work; (c) the change of enthalpy; (d) the change of internal energy.


1
Expert's answer
2021-11-21T17:30:31-0500

A 2-kg mass of oxygen expands at a constant-pressure of 172 kPa in piston-cylinder system from a temperature of 32°C to a final temperature of 182°C. [ Oxygen Cp = 0.918 kJ/kg · K]

(a.) Compute heat required in kJ

(b.) Compute steady-flow work in kJ

(c.) Compute total change in Total enthalpy in kJ

"Q= 272.8 \;kJ \\

W= -77.97 \;kJ \\

ΔH=272.8 \;kJ \\

n= \frac{2000}{32} = 62.5 \;mol \\

P=172 \;kPa \\

T_1 = 305 \;K \\

T_2 = 455 \;K \\

V_1 =\frac{nRT_1}{P} \\

= \frac{62.5 \times 8.314 \times 305}{172 \times 10^3} \\

= 0.9214 \;m^3 \\

\frac{V_2}{V_1} = \frac{T_2}{T_1} \\

\frac{V_2}{0.9214} = \frac{455}{305} \\

V_2 = 1.3746 \;m^3

Work = W = -P_{ext}(V_2- V_1) \\

= -172 \times 10^3(\frac{nRT_2}{P} - \frac{nRT_1}{P}) \\

= -nR(T_2 -T_1) \\

= -62.5 \times 8.314 (455- 305) \\

= -77943.75 \;J \\

ΔE = nc_vΔT \\

= 62.5 \times \frac{5}{2} \times 8.314 \times 150 \\

= 194.86 \;kJ"

According to the first law of thermodynamics.

"ΔE = q+W \\



194.86 = q -77.94 \;kJ \\



q = 272.8 \;kJ \\



ΔH = nc_pΔT \\



= 62.5 \times \frac{7}{2} \times 8.314 \times 150 \;J \\



= 272.8 \;kJ"